OpenStudy (anonymous):

a + 5c − b − c + a − b

OpenStudy (anonymous):

2(a+2c-b)

OpenStudy (anonymous):

thats the answer

jimthompson5910 (jim_thompson5910):

optionally you can distribute to get 2a + 4c - 2b and you get this by saying a+a = 2a -b-b = -2b 5c - c = 4c

OpenStudy (anonymous):

Just my opinion, i think my answer is right, so when you do get it wrong or right can you please tell me the answer

OpenStudy (anonymous):

I disagree with Jim :/

OpenStudy (anonymous):

a + 5c − b − c + a − b 2a+4c-2b 2(a+2c-b) I'm with Amanda

OpenStudy (anonymous):

:) thank you

jimthompson5910 (jim_thompson5910):

my version is the same as yours amandasoto144, I just didn't factor out the 2

jimthompson5910 (jim_thompson5910):

so either answer works

OpenStudy (anonymous):

hey, hey, im not that smart ok LOL :)

jimthompson5910 (jim_thompson5910):

yes you're smart, you just overlooked that fact, which is ok